OpenStudy (anonymous):

Line Integrals: Integrate F(x,y,z)=x+y^1/2-z^2 over the path from (0,0,0) to (1,1,1) C1UC2 C1 r(t)=ti+t^2j from 0 to 1 C2 r(t)=i+j+tk from 0 to 1 Cannot get the correct answer. Help me find my mistake.

OpenStudy (anonymous):

Did you parameterize the line segment correctly, using (1-t)<ro> +t<r1> . I'm pretty sure that is the first step. Line integrals confuse me. I'll definitely be hitting the books this weekend.

OpenStudy (anonymous):

Yes, did that. but their answer and mine are way off.

OpenStudy (anonymous):

I beleive the formula is int from a to b f(x,y,z)ds

OpenStudy (anonymous):

translates to int 0 to 1, (t,t^2,0)(2^1/2)dt

OpenStudy (anonymous):

+ int 0 to 1 f(1,1,t)(1)dt

OpenStudy (anonymous):

first integral I get 2^1/2

OpenStudy (anonymous):

second integral i get 5/3+2^1/2

OpenStudy (anonymous):

Don't take my word for it. but I think the first step would be to find the line integral of c1, which would be f(r(t) (dot)r'(t) dt. I think this is what it should look like \[\int\limits\limits_{0}^{0}1+t * \sqrt{1^2 +(2t)^2 dt}\] Possibly?

OpenStudy (anonymous):

is that not a vector field? instead of a line integral?

OpenStudy (anonymous):

err, I think you are right. aha :[ Maybe someone with more knowledge should answer this.

OpenStudy (anonymous):

thanks anyways, I'm going to press on.

OpenStudy (anonymous):

Did you get this? \[\frac{1}{6} \left(11+5 \sqrt{5}\right)\]